Can you please help me with this math problem? I am having a difficult time with it.
There is a .6 probabilty that an elem tree has Dutch elem disease. An arborist selects 3 elms in the area. Give the sample space that list all of the possible combinations of diseased (D) and not diseased (N) for the 3 trees.
a) what is the probabilty that none of the 3 trees have the Dutch elm disease?
b) what is the probabilty that at lease one of the 3 trees has the Dutch elm disease?

I really appreciate you help with this question. Thank you so much!

Answer by stanbon(75887)

There is a .6 probabilty that an elem tree has Dutch elem disease. An arborist selects 3 elms in the area.
----------
Give the sample space that list all of the possible combinations of diseased (D) and not diseased (N) for the 3 trees.
DDD,DDN,DND,NDD,NND,NDN,DNN,NNN
----------------------------------------------
a) what is the probabilty that none of the 3 trees have the Dutch elm disease?
P(NNN) = 1/8
------
b) what is the probabilty that at least one of the 3 trees has the Dutch elm disease?
P(at least one has) = 1 - P(none has) = 1 - (1/8)
================
Cheers,
Stan H.
================